OpenStudy (anonymous):

What is the slope of the line that contains the points (1,5) and (3,2)? Choice A -2/3 Choice B 3/2 Choice C 2/3 Choice D -3/2

OpenStudy (anonymous):

D

OpenStudy (amistre64):

the slope is the change in y --------- change in x (1,5) -(3,2) ------ -2,3 ; slope = y/x = -3/2